OpenStudy (anonymous):

the diameter of the larger circle is 10.5 cm, the diameter f the smaller circle is 4.5 cm. What is the approximate area of the shaded region?

OpenStudy (campbell_st):

find the radius of each circle... and the formula for area of the circle is \[A = \pi r^2\] the if the circles have the same centre, you can use Shaded Area = Area of the large circle - Area of the small circle. hope it helps And the answer posted above by @happy_to_help is incorrect... that why Open Study asks people not to post answers.

OpenStudy (happy_to_help):

Sorry that i put the answer but to be exact my answer was correct @campbell_st

OpenStudy (campbell_st):

@happy_to_help in your answer you squared the diameter... you need to find the radius 1st... if you want to use the diameter you would need to divide the answer by 4 \[A = \pi (\frac{d}{2})^2 ~~~or~~~~A = \pi \times \frac{d^2}{4}\]

OpenStudy (triciaal):

what @campbell_st is correct this is just visual L = large S= small A = area both circles have the same center